### v1

# EdgeAlpha: Bringing Process Discovery to the Data Sources

Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT). The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, a distributed algorithm for process discovery operating directly on sensor nodes and edge devices on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. EdgeAlpha enables (a) scalable mining, as a node, for each event, only interacts with its predecessors and, when queried, only exchanges aggregates, i.e., partial footprint matrices, with the central location and (b) privacy preserving process mining, as nodes only store their own as well as predecessor and successor events. On the Sepsis Cases event log, for example, a node queries on average 18.7% of all nodes. For the Hospital Log, we can even reduce the overall querying to 3.87% of the nodes.
Keywords:
Distributed Process Mining Event Data Stream In-Network Processing Scalability
1 Introduction
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from IoT sensors, for example, in a smart manufacturing or smart health setting [6, 13, 17]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [8, 15, 24]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [3]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place [14, 19, 21]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes events directly at the source (right), and only exchanges partial footprint matrices.
In this paper, we introduce EdgeAlpha, a distributed algorithm for process mining based on the Alpha Miner [4]. Utilizing the constrained compute and storage resources of sensor nodes and edge devices we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. For each event, that a node in EdgeAlpha senses, it locally stores the event as well as the predecessor and successor of this event and updates a local footprint matrix (FM) accordingly, see Figure 1. Thus, over time, each node constructs a partial FM that represents the partial view of this node on the process. To derive the complete process model, a central entity queries the nodes for their partial FM and combines these to a complete matrix.
To determine the predecessor of an event, the node that records this particular event, has to potentially query all other nodes, which can lead to large communication overheads. In this paper, we show that events of real-world processes statistically only have a few predecessors, which further only rarely change throughout the lifetime of a process. As a result, for each event, a node can very efficiently determine the predecessor, which in turn reduces the need for messaging within the network and thereby increases scalability. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner. EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [22].
Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [9] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. This observation enables EdgeAlpha to reduce the communication overhead for determining predecessor events by up to 96% compared to querying all nodes in the network. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. By batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor as well as a baseline that always queries the correct node immediately.
In summary, the contributions of this paper are as follows.
  • We introduce EdgeAlpha, a distributed Alpha Miner operating directly where the events originate.
  • We analytically and experimentally show the correctness of the algorithm.
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach when querying after each event.
  • We ensure that nodes only share aggregated event data with the central entity.
The remainder of the paper is structured as follows. Section 2 introduces the necessary background. Next, Section 3 presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section 4. In Section 5, we evaluate EdgeAlpha. We conclude the paper in Section 6.
2 Preliminaries and Background
In this section, we provide the required preliminaries and background. We begin with definitions of direct successions and footprint matrices, and then recapitulate the Alpha Miner Algorithm [4].
We refer to Peterson [20] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [1, 5]. In the following, let $L$ be an event log over the activity set $T$. We define direct successions and the resulting relationships between two activities following the definitions by Gatta et al. [14].
Definition 1 (Direct Succession)
Activity $b$ directly succeeds $a$ in event log $L$, i.e. $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{1},\dots,t_{n})\in L$ and $j\in\{1,\dots,n-1\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
Definition 2 (Causality, No Direct Succession)
With the help of Definition 1 we now define causality (1) and no direct succession (2) between two activities.
$a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (1)
$a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
Next, we recapitulate the fundamentals of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm, see Algorithm 1, identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Each activity of this set corresponds to a transition in the final output model computed by the algorithm. Next, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and the end activities of these traces, respectively. Subsequently, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets where each element of the first set is directly followed by every element of the second set, although not vice versa. Moreover, activities within each set do not directly succeed one another. The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step. Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
Algorithm 1 Alpha Miner
1:$T_{L}~{}=\{t\in T~{}|~{}\exists~{}\sigma\in L:~{}~{}t\in\sigma\}$
2:$T_{I}~{}=\{first(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
3:$T_{O}=\{last(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
[…]
6:$P_{L}=\{p_{(A,B)}~{}|~{}(A,B)\in Y_{L}\}~{}\cup~{}\{i_{L},o_{L}\}$
[…]
8:$\alpha(L)=(P_{L},~{}T_{L},~{}F_{L})$
As part of the calculation of $X_{L}$ in the Alpha Miner, it first computes a so-called footprint matrix (FM). Essentially, the FM serves to store information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
$FM_{ij}=1\quad$ $\Leftrightarrow\quad i>_{L}j$
$FM_{ij}=0\quad$ $\Leftrightarrow\quad i\not>_{L}j$
For our design, we additionally define partial footprint matrices that contain only one specific column. For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
3 Related Work
Van der Aalst [2] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [11] leverage Map-Reduce [7] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [12] extends their approach to the Flexible Heuristics Miner [23] and evaluates the performance of the Map-Reduce based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research places dual emphasis on scalability, i.e, processing data directly at the source, and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [14] address privacy concerns in the healthcare domain: Here, each hospital center aggregates their own event data into an FM, such that the others only see the aggregated versions of the data. Just as other works [2, 11, 12], they merge the FMs subsequently. Overall, all these collect event logs at one or a few locations. In EdgeAlpha, in contrast, each node collects a local event log and only shares aggregates, i.e., partial FMs, with the central location. As a result, EdgeAlpha increases scalability and reduces privacy risks.
4 Design
In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
4.1 Assumptions on Events and Event Logs
In adherence to common practice, we assume events with the attributes case identifier, activity name and timestamp, whereas the case identifier is unique for each case. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we – without loss of generality – assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
4.2 EdgeAlpha Algorithm
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing EdgeAlpha algorithm. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node to detect exactly one type of activity. Note, that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types. Furthermore, we assume that all nodes can communicate reliably and without significant delays, i.e., use common communication technologies and protocols such as cellular connectivity, Wi-Fi or Bluetooth.
EdgeAlpha as algorithm has two phases: Phase 1, initiated upon event triggering, and Phase 2, activated upon a process model request. In phase 1, whenever a node records an event, it finds the predecessor of the event by communicating to other nodes in the system and, once it determines the predecessor, it notes the directly follows relation in its local FM. As a result, every node continuously updates its local, partial FM over time. Phase 2 is activated whenever an entity requests a process model. Then all nodes send their partial FM to a central location which then constructs a FM from these partial ones. Overall, phase 2 is activated on demand and triggered by outside events, while phase 1 operates continuously. In the following, we define both phases. For simplicity, we begin by introducing the basic version of Phase 1, which, for each event, sends a message to each node in the system. Later, we introduce a more efficient variant and event batching.
4.2.1 Phase 1 – Event Ordering and Partial Footprint Matrix Construction.
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Thus, after the node senses the event, it queries for the predecessor event, to which it does not get any responses. Thus, it sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response including event 1. It increments the value in the corresponding cell of its local FM.
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps, see Figure 2 for an example. First, node $i$ stores event $e$ locally, including its case identifier, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp, that is still smaller than the own one, as predecessor. We denote the predecessor node as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
The case where there are multiple events to choose from, indicates communication delays. If delays result in incorrectly set predecessors or successors, the nodes rectify the stored events once the correct event is identified and update the counts in the successor’s partial FM accordingly. This is why nodes store a partial FM that not only encodes whether a direct succession has occurred, but also count the occurrences.
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessors in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it directly.
4.2.2 Phase 2 – Requesting Process Data.
This phase begins when an outside entity, for example a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Figure 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries that are greater than 0 to 1, which aggregates the data even more and pushes it back into the conventional FM structure. To determine whether its respective activity is an end activity, it verifies whether it detected an event lacking a successor. The requesting unit receives all responses, it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating all columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity identifier. Further, we merge the start and end activities into the sets $T_{I}$ and $T_{O}$, respectively, following the conventions of the Alpha Miner. Subsequently, the requesting unit proceeds with the remainder of the algorithm, starting from step four, the determination of $X_{L}$, which includes calculating the FM, here $FM^{all}$, see Algorithm 1.
Figure 3: Phase 2 – Requesting Process Data: First, we request the partial FMs and whether the activity was the start or end of any trace from all nodes. Upon receiving the information, we concatenate the matrices, form start and end activity sets, compute the process model.
4.2.3 Most-Frequent-Predecessors (MFPs).
In phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor as follows: we query nodes in order of their likeliness to be the predecessor. Thus, when searching for the predecessor of an event, a node first contacts the most likely predecessor and then the second most likely one and so on. For this, each node records how often each predecessor occurs. Given that activities are, in real-world scenarios, preceded by a consistent subset of activities, this approach prioritizes querying the most frequently occurring predecessors. When detecting a start event, the corresponding node has to still query every single node in the network, to be sure that it is the start event. In our evaluation in Section 5 we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures the scalability of EdgeAlpha.
4.2.4 Batching.
To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events. Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message. The nodes continue to use MFP Requesting. As soon as the node finds the predecessor of an event, it removes the event from the batch. The node ceases to query additional nodes once the batch is empty.
4.2.5 Correctness.
After introducing the algorithmic design of EdgeAlpha, we next prove its correctness. For this, we show in the following that EdgeAlpha outputs the same process models as the traditional centralized Alpha Miner algorithm.
We begin by introducing notations before defining our claim. Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes all activities in $L$. Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of. Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column. In the following correctness claim, we omit start and end activities. The design of EdgeAlpha ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace. By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
Claim
Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity identifier results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
Proof
To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa. Thus, proving the claim of equivalence between the merged partial FMs and the FM with the unrestricted knowledge.
By definition, $FM^{all}\in\{0,1\}^{n\times n}$. As already discussed, for $FM^{i},~{}i\in\{0,\dots,n-1\}$, we assume without loss of generality that $FM^{i}\in\{0,1\}^{n\times 1}$. It directly follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}\in\{0,1\}^{n\times n}$.
Let $i>_{L_{j}}j,~{}i,j\in\{0,\dots,n-1\}$ be a direct succession encoded in the merged partial footprint matrix $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}$. Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$. $FM^{j}$ contains all direct successions between its own activity and the activities preceding it. This means, there is a trace in the event log $L$ in which $i$ is followed by $j$. By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity identifier. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus, proving the claim.
$\Box$
4.2.6 Design Discussion.
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead, in EdgeAlpha, each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
Further, the design of EdgeAlpha is directly applicable to process mining algorithms beyond the Alpha Miner. Algorithms that start by building a FM or keep track of direct successions in the event log, can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [22]. The vast majority of results that we show in our evaluation in Section 5, such as the efficiency of Phase 1 of EdgeAlpha, is not specific to the Alpha Miner and directly applies, for example, to the Heuristics Miner.
5 Evaluation
In this section, we evaluate EdgeAlpha in terms of performance. We start by introducing our implementation in Section 5.1 and suitable baselines in Section 5.2. In Section 5.3, we introduce on which well-known datasets we conduct the experiments. We analyze these datasets in Section 5.4. In Section 5.5, we examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying of all nodes. We continue with an estimation on what FM sizes to expect during an algorithm run in Section 5.6. We finish the section by discussing the overall scalability of EdgeAlpha in Section 5.7
5.1 Implementation
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha. We also validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^1] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section 5.3.
5.2 Baselines
In our experiments, we compare EdgeAlpha with and without MFP Requesting. Not using MFP Requesting, i.e. querying all nodes (”Query All”), serves as upper bound. On the lower end, a theoretical, oracle-like, approach bounds the number of queried nodes in EdgeAlpha. It requires no requests if the predecessor corresponds to the node itself and otherwise takes exactly one request to identify the predecessor.
EdgeAlpha assumes a setting in which event detection is part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [4] or other distributed Alpha Miners such as the Map-Reduce approach [11] or the Alpha Miner in the healthcare scenario [14] is not viable. The novelty of EdgeAlpha lies in leveraging the event detection phase to pre-process events, a feature absent in other approaches.
5.3 Datasets
We conduct experiments on the datasets Hospital Log [9], Sepsis Cases [18] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [17]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [10] and the Road Traffic Fine Management Process [16]. Table 1 shows general statistics of the datasets, including the number of events, the number of cases, the average case length, and the number of activities. As each node is responsible for a single activity, the dataset, for example, represents IoT networks with size of 11 nodes in case of the Traffic Fine dataset and 624 nodes in case of the Hospital Log.
[…]
Traffic Fine 561,470 150,370 3.73 11
Sepsis 15,214 1,050 14.49 16
Smart Factories 8,607 271 31.76 21
BPIC 2017 1,202,267 31,509 38.16 26
Hospital 150,291 1,143 131.49 624
Table 1: Key data of the datasets showcasing their basic properties such as event counts, the number of cases, the average case length or the number of activities.
Event Log ##\##Start Activities ##\##Self-Loops Avg ##\##pred. (SD) Avg MFP Ratio (SD) Reduction
Traffic Fine 1 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
Sepsis 6 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
Smart Factories 1 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
BPIC 2017 1 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
Hospital 29 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
Table 2: Dataset details that directly influence the number of requests in phase 1 of EdgeAlpha. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
5.4 Dataset Analysis
We first analyze the properties of real-world event logs, see Table 2. For instance, our analysis shows that the Smart Factories event log contains numerous self-loops, which arise due to events having different lifecycle transitions and states in this event log. This characteristic is essential for EdgeAlpha, as self-loops do not require communication. Further, the table shows that activities commonly have only a few predecessors, underlining the potential of MFP Requesting.
5.5 Number of Queried Nodes
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node needs to query. In datasets like the Hospital Log there are up to 623 nodes to query, the network traffic can be considerably large, particularly during concurrent events. MFP Requesting in EdgeAlpha builds on past event data to optimize request in the future. For the first experiments we assume a batch size of one to focus on other characteristics. As we show in the the batch size analysis, the different batch sizes strongly influence the number of queried nodes.
Figure 4: Comparison of average queried nodes with MFP Requesting compared to querying all nodes. EdgeAlpha is able to reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
Figure 5: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. This diagram shows the distribution of queried node counts. A steeper curve, such as for the Hospital Log, indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
Figure 6: Relationship between batch size and average number of queried nodes per event across all datasets. To be able to compare the datasets with each other, we normalize the average number of queried nodes. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size the reduction rate decreases.
Figure 7: Stabilization of queried node count over time in Hospital Log. Since the data is fluctuating, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
Figure 8: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log. To put the averages into relation, we plot the number of requested nodes when querying all nodes as well as the average number of queried nodes over all events. The diagram indicates that the events that query on average the most nodes, occur the least. Start activities are excluded from this conclusion.
5.5.1 Comparison Across Datasets.
Figure 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table 2 also summarizes the significant reduction in the number of queried nodes. Notably, this reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
Figure 6 presents Cumulative Distribution Functions (CDF) comparing all datasets. The Hospital Log shows the steepest increase with over 90% of events requiring queries to no more than 5% of the nodes. Conversely, the Road Traffic Fine Management Process has barely any occurrences where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
5.5.2 Stabilization Analysis.
We now delve deeper into the specifics of the Hospital Log and the Sepsis Cases log. Figure 7 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log serves as a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
5.5.3 Activity-wise Analysis.
Figure 8 takes a closer look at the specifics of the Sepsis Cases event log. Depending on the activity under examination, the Sepsis Cases log has varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have a designated start event, but six different ones. However, in about 95% of the cases, a case starts with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count is the result of the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
5.5.4 Batch Size Analysis.
Figure 6 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log those are on average 15.6 nodes.
Overall, our results indicate that the reduction in the number of queried nodes depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors of each activity. The experiments show that it takes the activities some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a sliding window approach, where nodes track activity predecessors within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
5.6 Footprint Matrix Sizes
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating if a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in their FMs, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log that has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [11, 14], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
5.7 Evaluation Discussion
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [11, 14] which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
6 Conclusion
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored for process discovery in IoT networks. By decentralizing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how the design principles of EdgeAlpha can be applied to other process mining algorithms that construct footprint matrices or record direct successions in the beginning, such as the Heuristics Miner. Moreover, the integration of Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 50 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We highlight, how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.
{credits}
6.0.1 Acknowledgements
This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.

### v3

@@
-EdgeAlpha: Bringing Process Discovery to the Data Sources
+EdgeAlpha: Distributed Process Discovery at
 Abstract
-Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT).
+Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT) in a smart home, factory or hospital setting.
 The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs.
 Further, privacy considerations often prevent data collection at a central location in the first place.
-To address this challenge, this paper introduces EdgeAlpha, a distributed algorithm for process discovery operating directly on sensor nodes and edge devices on a stream of real-time event data.
+To address this challenge, this paper introduces EdgeAlpha, an algorithm for distributed process discovery operating directly on sensor nodes on a stream of real-time event data.
 Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded.
 From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model.
-EdgeAlpha enables (a) scalable mining, as a node, for each event, only interacts with its predecessors and, when queried, only exchanges aggregates, i.e., partial footprint matrices, with the central location and (b) privacy preserving process mining, as nodes only store their own as well as predecessor and successor events.
-On the Sepsis Cases event log, for example, a node queries on average 18.7% of all nodes.
-For the Hospital Log, we can even reduce the overall querying to 3.87% of the nodes.
-Keywords:
-Distributed Process Mining Event Data Stream In-Network Processing Scalability
-1 Introduction
-In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from IoT sensors, for example, in a smart manufacturing or smart health setting [6, 13, 17].
+Our analysis shows that even with over 600 activities, like in the Hospital Log, the average number of predecessors per activity remains below 7.
+This enables efficient determination of predecessors for each event, reducing the communication overhead by up to 96% compared to querying all nodes and enhancing scalability.
+Additionally, our analysis demonstrates that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups of 40 reduces the average queried nodes per event to less than 2.5%.
+EdgeAlpha ensures that the resulting process model is identical to the original Alpha Miner.
+Index Terms:
+Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability
+I Introduction
+In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from Internet of Things (IoT) sensors, for example, in a smart manufacturing, smart home or smart health setting [1, 2, 3, 4].
 However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [5, 6, 7].
 This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [8].
-Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place [14, 19, 21].
+Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place.
+For example, patient data in hospitals, personal data of workers, data in smart home setups, or company secrets in cross-organizational processes are commonly not shared for analysis [9, 10, 11].
 We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
-Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes events directly at the source (right), and only exchanges partial footprint matrices.
-In this paper, we introduce EdgeAlpha, a distributed algorithm for process mining based on the Alpha Miner [4].
-Utilizing the constrained compute and storage resources of sensor nodes and edge devices we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location.
-For each event, that a node in EdgeAlpha senses, it locally stores the event as well as the predecessor and successor of this event and updates a local footprint matrix (FM) accordingly, see Figure 1.
-Thus, over time, each node constructs a partial FM that represents the partial view of this node on the process.
-To derive the complete process model, a central entity queries the nodes for their partial FM and combines these to a complete matrix.
-To determine the predecessor of an event, the node that records this particular event, has to potentially query all other nodes, which can lead to large communication overheads.
-In this paper, we show that events of real-world processes statistically only have a few predecessors, which further only rarely change throughout the lifetime of a process.
-As a result, for each event, a node can very efficiently determine the predecessor, which in turn reduces the need for messaging within the network and thereby increases scalability.
-Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner.
+Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
+In this paper, we introduce EdgeAlpha, an algorithm for distributed process mining based on the Alpha Miner [12].
+Utilizing the constrained compute and storage resources of sensor nodes, further referred to as nodes, we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location.
+In contrast to Gatta et al.
+[9] which enable cross-organizational process mining through the transmission of data aggregates, EdgeAlpha distributes the computation of the footprint matrices (FMs) to the sensor-node level.
+This approach allows the collection of aggregated data within an organization that would otherwise not be obtainable in raw form due to, for example, too high data velocity or data protection regulations.
+For each event that a node in EdgeAlpha detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig.
+1.
+Thus, over time, each node constructs a partial FM that represents the view of this node on the process.
+To derive the complete process model, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix.
+EdgeAlpha enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with a central entity and (b) privacy-preserving process mining, as nodes only store their own as well as predecessor and successor events.
+To determine an event’s predecessor event, the node that records this particular event potentially has to query all other nodes, which can lead to large communication overheads.
+In this paper, we show that activities of real-world processes statistically only have a few predecessors activities, which further only rarely change throughout the lifetime of a process.
+Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [13] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors.
+As a result, for each event, a node can very efficiently determine the predecessor reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability.
+Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events.
+Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
+In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor and a baseline that immediately queries the correct node.
+Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner and show that in four out of five datasets an intermediate result FM already complies with the FM of the Alpha Miner by more than 90% after less than 5% of the events of the datasets.
 EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [14].
-Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [9] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors.
-This observation enables EdgeAlpha to reduce the communication overhead for determining predecessor events by up to 96% compared to querying all nodes in the network.
-Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events.
-By batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
-In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor as well as a baseline that always queries the correct node immediately.
 In summary, the contributions of this paper are as follows.
@@
   • We analytically and experimentally show the correctness of the algorithm.
-  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach when querying after each event.
-  • We ensure that nodes only share aggregated event data with the central entity.
+  • We demonstrate that EdgeAlpha’s intermediate results quickly converge to the final result of the Alpha Miner.
+  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach and by 97.5% when batching queries with a batch size of 40.
+  • We ensure that the nodes only share aggregated event data with a central entity.
+  • We implement EdgeAlpha and publish its code open-source [^1] Available at: redacted..
 The remainder of the paper is structured as follows.
-Section 2 introduces the necessary background.
-Next, Section 3 presents related work on distributed process discovery.
-Afterwards, we introduce the design of EdgeAlpha in Section 4.
-In Section 5, we evaluate EdgeAlpha.
-We conclude the paper in Section 6.
-2 Preliminaries and Background
-In this section, we provide the required preliminaries and background. We begin with definitions of direct successions and footprint matrices, and then recapitulate the Alpha Miner Algorithm [4].
+Section II introduces the necessary background.
+Next, Section III presents related work on distributed process discovery.
+Afterwards, we introduce the design of EdgeAlpha in Section IV.
+In Section V, we evaluate EdgeAlpha.
+We conclude the paper in Section VI.
+II Preliminaries and Background
+In this section, we provide the required preliminaries and background.
+We provide definitions of direct successions and footprint matrices and recapitulate the steps of the Alpha Miner Algorithm [12] that are relevant to our approach.
 We refer to Peterson [15] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [16, 17].
@@
 Activity $b$ directly succeeds $a$ in event log $L$, i.e.
-$a>_{L}b$, if and only if there exists a trace $\sigma=(t_{1},\dots,t_{n})\in L$ and $j\in\{1,\dots,n-1\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
+$a>_{L}b$, if and only if there exists a trace $\sigma=(t_{0},\dots,t_{n-1})\in L$ and $j\in\{0,\dots,n-2\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
 Definition 2 (Causality, No Direct Succession)
@@
 $a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
-Next, we recapitulate the fundamentals of the Alpha Miner before later in the paper introducing a distributed variant.
-First, the algorithm, see Algorithm 1, identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$.
-Each activity of this set corresponds to a transition in the final output model computed by the algorithm.
-Next, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and the end activities of these traces, respectively.
-Subsequently, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets where each element of the first set is directly followed by every element of the second set, although not vice versa.
-Moreover, activities within each set do not directly succeed one another.
-The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step.
-Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
-Algorithm 1 Alpha Miner
-1:$T_{L}~{}=\{t\in T~{}|~{}\exists~{}\sigma\in L:~{}~{}t\in\sigma\}$
-2:$T_{I}~{}=\{first(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
-3:$T_{O}=\{last(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
-[…]
-[…]
-6:$P_{L}=\{p_{(A,B)}~{}|~{}(A,B)\in Y_{L}\}~{}\cup~{}\{i_{L},o_{L}\}$
-[…]
-8:$\alpha(L)=(P_{L},~{}T_{L},~{}F_{L})$
-As part of the calculation of $X_{L}$ in the Alpha Miner, it first computes a so-called footprint matrix (FM).
-Essentially, the FM serves to store information regarding the direct succession relationships between activities within the event log.
-For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is element of $\{0,\dots,|T_{L}|-1\}$.
+Next, we recapitulate the initial steps of the Alpha Miner before later in the paper introducing a distributed variant.
+First, the algorithm identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$.
+Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively.
+Subsequently, the algorithm determines all direct successions in $L$ which can be stored in an FM.
+This FM is then the basis for the further calculations in the algorithm and the resulting Petri net.
+A FM stores information regarding the direct succession relationships between activities within the event log.
+For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$.
 This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
@@
 $FM_{ij}=0\quad$ $\Leftrightarrow\quad i\not>_{L}j$
-For our design, we additionally define partial footprint matrices that contain only one specific column.
-For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
-3 Related Work
+For our design, we define partial footprint matrices containing only one specific column: For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
+III Related Work
 Van der Aalst [18] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log.
@@
 However, their approach still relies on an event log that is initially stored at a central location.
-Subsequent work [12] extends their approach to the Flexible Heuristics Miner [23] and evaluates the performance of the Map-Reduce based Alpha Miner and the Flexible Heuristics Miner.
-While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research places dual emphasis on scalability, i.e, processing data directly at the source, and privacy preservation by only storing data aggregates at a central location.
+Subsequent work [21] extends their approach to the Flexible Heuristics Miner [22] and evaluates the performance of the Map-Reduce-based Alpha Miner and the Flexible Heuristics Miner.
+While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only storing data aggregates at a central location.
 Gatta et al.
-[14] address privacy concerns in the healthcare domain: Here, each hospital center aggregates their own event data into an FM, such that the others only see the aggregated versions of the data.
-Just as other works [2, 11, 12], they merge the FMs subsequently.
-Overall, all these collect event logs at one or a few locations.
-In EdgeAlpha, in contrast, each node collects a local event log and only shares aggregates, i.e., partial FMs, with the central location.
-As a result, EdgeAlpha increases scalability and reduces privacy risks.
-4 Design
+[9] address privacy concerns in the healthcare domain: In their work, each hospital center aggregates its own event data into an FM, such that the others only see the aggregated versions of the data.
+Like other works [18, 19, 21], they subsequently merge the FMs.
+In contrast, EdgeAlpha distributes the FM computations to the sensor nodes that record the events, enabling data aggregation at an earlier stage.
+This allows organizations to collect and utilize data in aggregated form that would otherwise be restricted in its raw form due to data protection regulations.
+As a result of processing events directly at their sources, EdgeAlpha increases scalability and reduces privacy risks.
+IV Design
 In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
-4.1 Assumptions on Events and Event Logs
-In adherence to common practice, we assume events with the attributes case identifier, activity name and timestamp, whereas the case identifier is unique for each case.
+IV-A Assumptions on Events and Event Logs
+In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case.
+Our evaluation shows these assumptions are both realistic and applicable in numerous real-world scenarios.
+For example, in a hospital each patient is commonly assigned a fixed ID that can serve as a case ID and the events such as blood tests or ultrasound are predefined activities.
 Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log.
 Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log.
-As we operate on distributed event logs, we – without loss of generality – assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
-4.2 EdgeAlpha Algorithm
-First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing EdgeAlpha algorithm.
-Without loss of generality, and to ease simplicity and comprehensibility, we assume each node to detect exactly one type of activity.
-Note, that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
-Furthermore, we assume that all nodes can communicate reliably and without significant delays, i.e., use common communication technologies and protocols such as cellular connectivity, Wi-Fi or Bluetooth.
-EdgeAlpha as algorithm has two phases: Phase 1, initiated upon event triggering, and Phase 2, activated upon a process model request.
-In phase 1, whenever a node records an event, it finds the predecessor of the event by communicating to other nodes in the system and, once it determines the predecessor, it notes the directly follows relation in its local FM.
-As a result, every node continuously updates its local, partial FM over time.
-Phase 2 is activated whenever an entity requests a process model.
-Then all nodes send their partial FM to a central location which then constructs a FM from these partial ones.
-Overall, phase 2 is activated on demand and triggered by outside events, while phase 1 operates continuously.
-In the following, we define both phases.
-For simplicity, we begin by introducing the basic version of Phase 1, which, for each event, sends a message to each node in the system.
-Later, we introduce a more efficient variant and event batching.
-4.2.1 Phase 1 – Event Ordering and Partial Footprint Matrix Construction.
+As we operate on distributed event logs, we assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
+IV-B EdgeAlpha Setting
+First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing the EdgeAlpha algorithm.
+We employ EdgeAlpha in typical IoT application settings: A network of sensor nodes that are, for example, integrated into machines in a smart factory or smart health setting.
+These sensors can communicate with each other, either directly or via the Internet.
+Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications.
+Together, these allow a node to find the processor and successor of an event.
+Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP [23].
+This is essential, as accurate timestamps ensure proper event ordering.
+Without loss of generality, and to ease simplicity and comprehensibility, we assume each node detects exactly one type of activity.
+Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
+IV-C EdgeAlpha Algorithm: Overview
+EdgeAlpha consists of two phases: Phase 1 is initiated upon event triggering, and Phase 2 is activated upon a process model request.
+In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs.
+As a result, every node continuously updates its local, partial FM.
+Other than Phase 1, Phase 2 is activated on demand whenever an entity requests a process model.
+Then, all nodes send their partial FMs to a central location, which constructs an FM.
+We define both phases, starting with the basic version of Phase 1, which sends a message to all nodes in the system for each event, followed by a more efficient variant and event batching.
+IV-D Phase 1 – Event Ordering & Partial Footprint Matrices
 Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing.
 In this specific setting, event 1 is a start event.
-Thus, after the node senses the event, it queries for the predecessor event, to which it does not get any responses.
-Thus, it sets a flag that it detected a start event.
-For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response including event 1.
+Therefore, after the node detects the event, it queries for the predecessor event, for which it does not get any responses.
+Thus, the node sets a flag that it detected a start event.
+For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response, including event 1.
 It increments the value in the corresponding cell of its local FM.
-When an event $e$ on node $i$ is triggered, node $i$ executes the following steps, see Figure 2 for an example.
-First, node $i$ stores event $e$ locally, including its case identifier, the activity name and timestamp.
+When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig.
+2 for an example.
+First, node $i$ stores event $e$ locally, including its case ID, the activity name and timestamp.
 Then, it sends all other nodes a message containing $e$ to determine the predecessor event.
@@
 If such an event exists, the node responds with the potential predecessor event data.
-If there are multiple responses, node $i$ picks the event with the latest timestamp, that is still smaller than the own one, as predecessor.
-We denote the predecessor node as node $j$.
+If there are multiple responses, node $i$ picks the event with the latest timestamp that is still smaller than the own one, as predecessor.
+We denote the node with the predecessor event as node $j$.
 Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$.
 Subsequently, node $j$ saves $e$ as the successor event of its own event.
-The case where there are multiple events to choose from, indicates communication delays.
-If delays result in incorrectly set predecessors or successors, the nodes rectify the stored events once the correct event is identified and update the counts in the successor’s partial FM accordingly.
-This is why nodes store a partial FM that not only encodes whether a direct succession has occurred, but also count the occurrences.
-Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessors in its FM.
+If a node has to choose its predecessor from several events, it indicates communication delays which may cause incorrectly set predecessors.
+However, this is automatically rectified during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor.
+The node sets the correct successor and informs the node corresponding to the incorrect successor.
+This one starts the predecessor search again.
+The node rectifies the stored events once the correct event is identified and updates the counts in its partial FM accordingly.
+For this reason nodes store a partial FM that not only encodes whether a direct succession has occurred but also counts the occurrences.
+The rectification process works due to a single start event and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests.
+As soon as a node has more than one successor in a single trace the incorrectly set one is informed and an total order of events is established.
+Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM.
 Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities.
@@
 If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event.
-A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it directly.
-4.2.2 Phase 2 – Requesting Process Data.
-This phase begins when an outside entity, for example a central node, requests a process model.
-During this phase, we request the local FMs of all nodes and start or end events within any case, see Figure 3.
+A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
+EdgeAlpha’s design further incorporates a sliding-window-principle for discarding old data in order to respond to the requirements of the resource-restricted sensor nodes.
+With this principle, as soon as a certain memory size has been reached or a certain time has been exceeded, the node discards the oldest data when adding new data.
+However, the sliding window approach only gets beneficial with a vast amount of data.
+The event logs we consider in the evaluation do not fill enough memory to utilize the sliding window.
+IV-E Phase 2 – Requesting Process Data
+This phase begins when an outside entity, for example, a central node, requests a process model.
+During this phase, we request the local FMs of all nodes and start or end events within any case, see Fig.
+3.
 Subsequently, each node responds with its local FM.
-Before sending, the node changes all entries that are greater than 0 to 1, which aggregates the data even more and pushes it back into the conventional FM structure.
-To determine whether its respective activity is an end activity, it verifies whether it detected an event lacking a successor.
-The requesting unit receives all responses, it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating all columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity identifier.
-Further, we merge the start and end activities into the sets $T_{I}$ and $T_{O}$, respectively, following the conventions of the Alpha Miner.
-Subsequently, the requesting unit proceeds with the remainder of the algorithm, starting from step four, the determination of $X_{L}$, which includes calculating the FM, here $FM^{all}$, see Algorithm 1.
-Figure 3: Phase 2 – Requesting Process Data: First, we request the partial FMs and whether the activity was the start or end of any trace from all nodes.
-Upon receiving the information, we concatenate the matrices, form start and end activity sets, compute the process model.
-4.2.3 Most-Frequent-Predecessors (MFPs).
+Before sending, the node changes all entries greater than 0 to 1, aggregating the data even more.
+This converts the local partial FM back into the conventional FM structure.
+Each node checks if events without a successor were recorded to determine whether its respective activity is an end activity.
+The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID.
+Further, we merge the start and end activities into two sets $T_{I}$ and $T_{O}$, following the conventions of the Alpha Miner.
+Next, the requesting unit proceeds with the remainder of the Alpha Miner algorithm, starting after the calculation of the FM, here $FM^{all}$.
+Figure 3: Phase 2 – Requesting Process Data: We request partial FMs and start/end activity flags from all nodes.
+Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the process model.
+IV-F Most-Frequent-Predecessors (MFPs)
 In Phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log.
-To reduce communication overhead, we search for the predecessor as follows: we query nodes in order of their likeliness to be the predecessor.
-Thus, when searching for the predecessor of an event, a node first contacts the most likely predecessor and then the second most likely one and so on.
-For this, each node records how often each predecessor occurs.
-Given that activities are, in real-world scenarios, preceded by a consistent subset of activities, this approach prioritizes querying the most frequently occurring predecessors.
-When detecting a start event, the corresponding node has to still query every single node in the network, to be sure that it is the start event.
-In our evaluation in Section 5 we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures the scalability of EdgeAlpha.
-4.2.4 Batching.
+To reduce communication overhead, we search for the predecessor by querying nodes in order of their likelihood of being the predecessor.
+Thus, when searching for an event’s predecessor, a node first contacts the most likely predecessor, then the second most likely one, and so on.
+For this, during runtime each node continuously records how often each predecessor occurs.
+As we show later in a dataset analysis, most activities in real-world scenarios are preceded by a typically stable and small subset of activities, this approach prioritizes querying the most frequently occurring predecessors.
+However, when detecting a start event, the corresponding node still has to query every node in the network to ensure it is the start event.
+In our evaluation in Section V we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures scalability.
+IV-G Batching
 To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events.
@@
 The node ceases to query additional nodes once the batch is empty.
-4.2.5 Correctness.
+IV-H Correctness
 After introducing the algorithmic design of EdgeAlpha, we next prove its correctness.
@@
 By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
-Claim
-Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity identifier results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
-Proof
+Claim 1
+Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity ID results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
+Proof 1
 To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa.
-Thus, proving the claim of equivalence between the merged partial FMs and the FM with the unrestricted knowledge.
+Thus proving the claim of equivalence between the merged partial FMs and the FM with unrestricted knowledge.
 By definition, $FM^{all}\in\{0,1\}^{n\times n}$.
@@
 Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$.
-$FM^{j}$ contains all direct successions between its own activity and the activities preceding it.
-This means, there is a trace in the event log $L$ in which $i$ is followed by $j$.
+$FM^{j}$ contains all direct successions between its activity and its preceding activities.
+This means there is a trace in the event log $L$ in which $i$ is followed by $j$.
 By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
@@
 Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$.
-Merging all partial FMs only means concatenating them ordered by activity identifier.
+Merging all partial FMs only means concatenating them ordered by activity ID.
 Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$.
-Thus, proving the claim.
+Thus proving the claim.
 $\Box$
-4.2.6 Design Discussion.
+IV-I Design Discussion
 A major advantage of EdgeAlpha is that event logs are no longer stored in one central place.
-Instead, in EdgeAlpha, each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity.
+Instead each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity.
 In addition, we use batching, i.e.
 a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
-Further, the design of EdgeAlpha is directly applicable to process mining algorithms beyond the Alpha Miner.
-Algorithms that start by building a FM or keep track of direct successions in the event log, can be projected onto our distribution approach.
+Further, the design of EdgeAlpha directly applies to process mining algorithms beyond the Alpha Miner.
+Algorithms that start by building an FM or keeping track of direct successions in the event log can be projected onto our distribution approach.
 A well-known example of another algorithm is the Heuristics Miner [14].
-The vast majority of results that we show in our evaluation in Section 5, such as the efficiency of Phase 1 of EdgeAlpha, is not specific to the Alpha Miner and directly applies, for example, to the Heuristics Miner.
-5 Evaluation
-In this section, we evaluate EdgeAlpha in terms of performance.
-We start by introducing our implementation in Section 5.1 and suitable baselines in Section 5.2.
-In Section 5.3, we introduce on which well-known datasets we conduct the experiments.
-We analyze these datasets in Section 5.4.
-In Section 5.5, we examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying of all nodes.
-We continue with an estimation on what FM sizes to expect during an algorithm run in Section 5.6.
-We finish the section by discussing the overall scalability of EdgeAlpha in Section 5.7
-5.1 Implementation
+The vast majority of results that we show in our evaluation in Section V, such as the efficiency of Phase 1 of EdgeAlpha, are not specific to the Alpha Miner and directly apply, for example, to the Heuristics Miner.
+V Evaluation
+In this section, we evaluate EdgeAlpha in terms of correctness, accuracy, and performance.
+We present our implementation in Section V-A and suitable baselines in Section V-B.
+Section V-C introduces on which well-known datasets we conduct the experiments.
+We analyze these datasets in Section V-D.
+In Section V-E we discuss correctness experiments and the compliance between the FMs of EdgeAlpha and the traditional Alpha Miner.
+We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section V-F.
+Section V-G estimates FM sizes during an algorithm run and Section V-H discusses the overall scalability of EdgeAlpha.
+V-A Implementation
 We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha.
-We also validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^1] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section 5.3.
-5.2 Baselines
+V-B Baselines
 In our experiments, we compare EdgeAlpha with and without MFP Requesting.
-Not using MFP Requesting, i.e.
-querying all nodes (”Query All”), serves as upper bound.
-On the lower end, a theoretical, oracle-like, approach bounds the number of queried nodes in EdgeAlpha.
-It requires no requests if the predecessor corresponds to the node itself and otherwise takes exactly one request to identify the predecessor.
-EdgeAlpha assumes a setting in which event detection is part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning.
+”Query All” represents the upper bound where all nodes are queried, while a theoretical oracle-like approach serves as the lower bound.
+The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
+EdgeAlpha’s design assumes event detection as part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning.
 Thus, comparing EdgeAlpha to the traditional Alpha Miner [12] or other distributed Alpha Miners such as the Map-Reduce approach [19] or the Alpha Miner in the healthcare scenario [9] is not viable.
-The novelty of EdgeAlpha lies in leveraging the event detection phase to pre-process events, a feature absent in other approaches.
-5.3 Datasets
+The novelty of EdgeAlpha lies in utilizing the event detection phase to pre-process events, a feature absent in other approaches.
+V-C Datasets
 We conduct experiments on the datasets Hospital Log [13], Sepsis Cases [24] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version.
@@
 To further substantiate our evaluation, we also include the BPI Challenge 2017 [25] and the Road Traffic Fine Management Process [26].
-Table 1 shows general statistics of the datasets, including the number of events, the number of cases, the average case length, and the number of activities.
-As each node is responsible for a single activity, the dataset, for example, represents IoT networks with size of 11 nodes in case of the Traffic Fine dataset and 624 nodes in case of the Hospital Log.
+In our setting, each activity in a dataset represents a node in an IoT network.
+TABLE I: Dataset Characteristics (Standard deviation in parentheses if not stated otherwise).
 […]
-Traffic Fine 561,470 150,370 3.73 11
-Sepsis 15,214 1,050 14.49 16
-Smart Factories 8,607 271 31.76 21
-BPIC 2017 1,202,267 31,509 38.16 26
-Hospital 150,291 1,143 131.49 624
-Table 1: Key data of the datasets showcasing their basic properties such as event counts, the number of cases, the average case length or the number of activities.
-Event Log ##\##Start Activities ##\##Self-Loops Avg ##\##pred. (SD) Avg MFP Ratio (SD) Reduction
-Traffic Fine 1 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
-Sepsis 6 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
-Smart Factories 1 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
-BPIC 2017 1 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
-Hospital 29 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
-Table 2: Dataset details that directly influence the number of requests in phase 1 of EdgeAlpha.
+Traffic Fine 561,470 11 (1) 150,370 3.73 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
+Sepsis 15,214 16 (6) 1,050 14.49 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
+Smart Factories 8,607 21 (1) 271 31.76 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
+BPIC 2017 1,202,267 26 (1) 31,509 38.16 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
+Hospital 150,291 624 (29) 1,143 131.49 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
+V-D Dataset Analysis
+We analyze the properties of real-world event logs.
+Next to basic information such as the numbers of events, activities and cases and the average case length, Table I includes data that directly influences the number of requests in Phase 1.
 These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity.
 Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
-5.4 Dataset Analysis
-We first analyze the properties of real-world event logs, see Table 2.
-For instance, our analysis shows that the Smart Factories event log contains numerous self-loops, which arise due to events having different lifecycle transitions and states in this event log.
-This characteristic is essential for EdgeAlpha, as self-loops do not require communication.
-Further, the table shows that activities commonly have only a few predecessors, underlining the potential of MFP Requesting.
-5.5 Number of Queried Nodes
+Even in a dataset like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event.
+The dataset analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events.
+This characteristic benefits EdgeAlpha, as self-loops do not require communication.
+For self-loops, too, only in case of communication delays a self-loop may be set wrongfully for which, however, EdgeAlpha provides a rectification procedure.
+V-E Experimental Correctness and Fitness Analysis
+As a first step, we validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^3] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section V-C.
+Next, we examine how the merged FMs obtained by querying the nodes for their partial FMs after an increasing number of events differ from the intermediate FM produced by the traditional Alpha Miner.
+We compare them by using the footprint-based fitness value implemented in pm4py.
+Fig.
+7 confirms the full compliance of EdgeAlpha’s resulting FM with the Alpha Miner’s FM across all dataset.
+It also demonstrates that nearly all datasets reach 90% fitness with less than 5% of events.
+Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, although it is considerably small, as shown in Table I.
+The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events there are typically present in a common event log before requesting the first process model.
+Moreover, our analysis suggests that reducing the sliding window for discarding old data would have minimal impact on the resulting process model, but is dependent on the current dataset.
+V-F Number of Queried Nodes
 To assess EdgeAlpha’s performance, we analyze its two phases separately.
-The event ordering phase depends highly on the number of potential predecessors a node needs to query.
-In datasets like the Hospital Log there are up to 623 nodes to query, the network traffic can be considerably large, particularly during concurrent events.
-MFP Requesting in EdgeAlpha builds on past event data to optimize request in the future.
-For the first experiments we assume a batch size of one to focus on other characteristics.
-As we show in the the batch size analysis, the different batch sizes strongly influence the number of queried nodes.
-Figure 4: Comparison of average queried nodes with MFP Requesting compared to querying all nodes.
-EdgeAlpha is able to reduce communication by querying the most frequent predecessors.
+The event ordering phase depends highly on the number of potential predecessors a node queries, leading to significant network traffic, particularly during concurrent events, in datasets like the Hospital Log, where up to 623 nodes are queried.
+MFP Requesting in EdgeAlpha builds on past event data to optimize requests in the future.
+For the first experiments, we assume a batch size of one to focus on other characteristics.
+As our batch size analysis shows, the different batch sizes strongly influence the number of queried nodes.
+Figure 4: Comparison of average queried nodes with MFP Requesting and querying all nodes.
+Diagram also shows standard error.
+EdgeAlpha can reduce communication by querying the most frequent predecessors.
 With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
+Figure 5: Fitness over time.
+Intermediate result FMs of the data sets quickly converge into the FM of the traditional Alpha Miner.
+The BPIC 2017 data set, for example, already has a fitness value of over 90% after 200 events compared with the Alpha Miner FM, which is approx.
+0.02% of the events in the event log.
 Figure 6: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence.
-This diagram shows the distribution of queried node counts.
-A steeper curve, such as for the Hospital Log, indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
-Figure 6: Relationship between batch size and average number of queried nodes per event across all datasets.
-To be able to compare the datasets with each other, we normalize the average number of queried nodes.
+A steeper curve indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
+Figure 7: Relationship between batch size and average number of queried nodes per event.
+Normalizing the average queried nodes for comparison.
 Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50.
-With increasing batch size the reduction rate decreases.
+With increasing batch size, the reduction rate decreases.
 Figure 8: Stabilization of queried node count over time in Hospital Log.
-Since the data is fluctuating, we use moving averages to visualize the stabilization.
+Since the data fluctuates, we use moving averages to visualize the stabilization.
 A reference line (Query All) indicates the maximum number of requested nodes per event.
@@
 Figure 9: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis).
-The x-axis displays the different activities of the event log.
-To put the averages into relation, we plot the number of requested nodes when querying all nodes as well as the average number of queried nodes over all events.
-The diagram indicates that the events that query on average the most nodes, occur the least.
-Start activities are excluded from this conclusion.
-5.5.1 Comparison Across Datasets.
-Figure 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor.
+The x-axis displays the different activities of the event log and highlights the start activities (S).
+We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all event.
+The diagram shows that activities with the highest average queries occur the least, excluding start activities.
+V-F1 Comparison Across Datasets
+Fig.
+4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor.
 Query All denotes the naive way to message every node in the network.
-This serves as baseline to compare to.
+This serves as a baseline to compare to.
 With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event.
 We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant.
-Table 2 also summarizes the significant reduction in the number of queried nodes.
-Notably, this reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
-Figure 6 presents Cumulative Distribution Functions (CDF) comparing all datasets.
-The Hospital Log shows the steepest increase with over 90% of events requiring queries to no more than 5% of the nodes.
-Conversely, the Road Traffic Fine Management Process has barely any occurrences where the predecessor can be identified after querying just 5% of the nodes.
+Table I shows the significant reduction in the number of queried nodes.
+This reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
+Fig.
+7 shows Cumulative Distribution Functions (CDF) for all datasets.
+The Hospital Log has the steepest increase with over 90% of events needing queries to no more than 5% of the nodes.
+Conversely, the Traffic Fine dataset has few instances where the predecessor can be identified after querying just 5% of the nodes.
 This deviation stems from the limited number of 11 activities in the Traffic Fine dataset.
 Hence, within the 5% threshold, only self-loops are considered.
-5.5.2 Stabilization Analysis.
-We now delve deeper into the specifics of the Hospital Log and the Sepsis Cases log.
-Figure 7 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events.
+V-F2 Stabilization Analysis
+We now delve deeper into the specifics of the Hospital Log.
+Fig.
+8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events.
 There is still a higher amplitude after 5,000 events.
 However, there are no long-lasting deviations.
-This event log serves as a good example of stabilization due to its large activity set.
+This event log is a good example of stabilization due to its large activity set.
 Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
-5.5.3 Activity-wise Analysis.
-Figure 8 takes a closer look at the specifics of the Sepsis Cases event log.
-Depending on the activity under examination, the Sepsis Cases log has varied querying patterns.
-In contrast to other event logs, the Sepsis Cases do not have a designated start event, but six different ones.
-However, in about 95% of the cases, a case starts with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others.
+V-F3 Activity-wise Analysis
+Fig.
+9 takes a closer look at the specifics of the Sepsis Cases log.
+Depending on the activity, the Sepsis Cases log shows varied querying patterns.
+In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones.
+However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others.
 Although ”Release E” shows an increased query count, it is not a start activity.
-The increased query count is the result of the limited occurrence (only six times), indicating ongoing stabilization.
+The increased query count results from the limited occurrence (only six times), indicating ongoing stabilization.
 The situation is similar for the four activities to its left.
 They occur infrequently, which also explains the higher standard error.
-5.5.4 Batch Size Analysis.
-Figure 6 illustrates the reduction in the average number of queried nodes per event as the batch size increases.
+V-F4 Batch Size Analysis
+Fig.
+7 illustrates the reduction in the average number of queried nodes per event as the batch size increases.
 The rate in which the number of queries decreases varies between the datasets.
 However, for all datasets the rate decreases with a growing batch size.
-For all datasets, the average number of queried nodes per event already lies under 2.5% nodes in the network with a batch size of 40.
-For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log those are on average 15.6 nodes.
-Overall, our results indicate that the reduction in the number of queried nodes depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors of each activity.
-The experiments show that it takes the activities some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes.
+For all datasets, the average number of queried nodes per event already lies under 2.5% of nodes in the network with a batch size of 40.
+For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
+Overall, our results indicate that the reduction in the number of queries depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors per activity.
+The experiments show that it takes the nodes some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes.
 While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes.
-One such adaptation involves employing a sliding window approach, where nodes track activity predecessors within a fixed window.
+One such adaptation involves employing a smaller sliding window, where nodes track activity predecessors only within a fixed window.
 This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
-5.6 Footprint Matrix Sizes
+V-G Footprint Matrix Sizes
 In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages.
-These messages include the node’s local FM along with two additional Booleans indicating if a node recorded the start or end activity of a trace.
+These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace.
 Thus, the scalability of these messages depends highly on the size of each partial FM.
-In EdgeAlpha, each node only keeps track of their predecessors in their FMs, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities.
-Even for datasets like the Hospital Log that has an activity count of 624, each node transmits only a list consisting of at most 624 elements.
+In EdgeAlpha, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities.
+Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements.
 Compared to existing distributed Alpha Miners [19, 9], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size.
@@
 For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
-5.7 Evaluation Discussion
+V-H Evaluation Discussion
 EdgeAlpha leverages data locality to enhance its efficiency.
-Unlike existing distributed Alpha Miners [11, 14] which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source.
+Unlike existing distributed Alpha Miners [19, 9], which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source.
 As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm.
@@
 This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
-6 Conclusion
-In this paper, we introduce EdgeAlpha, a novel distributed approach tailored for process discovery in IoT networks.
-By decentralizing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches.
-We discuss how the design principles of EdgeAlpha can be applied to other process mining algorithms that construct footprint matrices or record direct successions in the beginning, such as the Heuristics Miner.
-Moreover, the integration of Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network.
+VI Conclusion
+In this paper, we introduce EdgeAlpha, a novel distributed approach tailored to process discovery in IoT networks.
+By distributing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches.
+We discuss how EdgeAlpha’s design principles can be applied to other process mining algorithms that initially construct footprint matrices or record direct successions, such as the Heuristics Miner [14].
+Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network.
 The batching of predecessor queries further minimizes communication overhead.
@@
 EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event.
-By batching predecessor queries in groups of 50 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes.
-We highlight, how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities.
+By batching predecessor queries in groups of 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes.
+We analytically and experimentally demonstrate the correctness of the algorithm and additionally show how the intermediate results of EdgeAlpha quickly converge to the final result of the Alpha Miner.
+We highlight how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same results as the Alpha Miner.
 By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.
-{credits}
-6.0.1 Acknowledgements
-This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.

### v4

@@
-EdgeAlpha: Distributed Process Discovery at
-Abstract
-Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT) in a smart home, factory or hospital setting.
-The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs.
-Further, privacy considerations often prevent data collection at a central location in the first place.
-To address this challenge, this paper introduces EdgeAlpha, an algorithm for distributed process discovery operating directly on sensor nodes on a stream of real-time event data.
-Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded.
-From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model.
-Our analysis shows that even with over 600 activities, like in the Hospital Log, the average number of predecessors per activity remains below 7.
-This enables efficient determination of predecessors for each event, reducing the communication overhead by up to 96% compared to querying all nodes and enhancing scalability.
-Additionally, our analysis demonstrates that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups of 40 reduces the average queried nodes per event to less than 2.5%.
-EdgeAlpha ensures that the resulting process model is identical to the original Alpha Miner.
-Index Terms:
-Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability
-I Introduction
-In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from Internet of Things (IoT) sensors, for example, in a smart manufacturing, smart home or smart health setting [1, 2, 3, 4].
-However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [5, 6, 7].
-This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [8].
-Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place.
-For example, patient data in hospitals, personal data of workers, data in smart home setups, or company secrets in cross-organizational processes are commonly not shared for analysis [9, 10, 11].
-We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
-Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
-In this paper, we introduce EdgeAlpha, an algorithm for distributed process mining based on the Alpha Miner [12].
-Utilizing the constrained compute and storage resources of sensor nodes, further referred to as nodes, we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location.
-In contrast to Gatta et al.
-[9] which enable cross-organizational process mining through the transmission of data aggregates, EdgeAlpha distributes the computation of the footprint matrices (FMs) to the sensor-node level.
-This approach allows the collection of aggregated data within an organization that would otherwise not be obtainable in raw form due to, for example, too high data velocity or data protection regulations.
-For each event that a node in EdgeAlpha detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig.
+EdgeMiner: Distributed Process Mining at the Data Sources
+Abstract.
+Process mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT).
+The volume and velocity of data generated by such sensors makes it increasingly challenging to efficiently process the data by traditional process discovery algorithms, which operate on a centralized event log.
+This paper presents EdgeMiner, an algorithm for distributed process mining operating directly on sensor nodes on a stream of real-time event data.
+In contrast to centralized algorithms, EdgeMiner tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded.
+As EdgeMiner aggregates direct successions on the individual nodes, the raw data does not need to be stored centrally, thus improving both scalability and privacy.
+We analytically and experimentally show the correctness of EdgeMiner.
+In addition, our evaluation results show that EdgeMiner determines predecessors for each event efficiently, reducing the communication overhead by up to 96% compared to querying all nodes.
+Further, we show that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups reduces the average queried nodes per event to less than 2.5%.
+Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability, Internet of Things
+1. Introduction
+Mining the never-ending data stream of events sensed by the Internet of Things (IoT) provides deep insights into, for example, processes in smart manufacturing, smart homes, and healthcare and is known as process mining (Bertrand et al., 2022; Fernandez-Llatas et al., 2015; Malburg et al., 2022; Janssen et al., 2020).
+However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud (Ge et al., 2018; Zaslavsky et al., 2013; van der Aalst, 2013).
+This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities (van der Aalst, 2016).
+Parsing such event logs, established algorithms commonly construct so-called footprint matrices (FMs), in which they store the order of events, so-called directly follows relations (van der Aalst et al., 2004; Leemans et al., 2013; Weijters et al., 2006).
+Next to the scalability challenge, such collection of events at a central location causes privacy concerns (Gatta et al., 2019; Mannhardt et al., 2019; Pika et al., 2020).
+Figure 1.
+While traditional process mining (left) collects all events at a central entity, EdgeMiner (right) processes them directly at the source, and only exchanges aggregates (partial footprint matrixes), increasing scalability and privacy.
+[…]
+[From centralized to distributed process mining.]While classical process mining (left) collects all events on a central entity, EdgeMiner processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
+In this paper, we introduce EdgeMiner, an algorithm for distributed process mining: it composes FMs, the key building-block of process mining, in a distributed manner within a network of sensor nodes.
+For each event that a node in EdgeMiner detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig.
 1.
-Thus, over time, each node constructs a partial FM that represents the view of this node on the process.
-To derive the complete process model, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix.
-EdgeAlpha enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with a central entity and (b) privacy-preserving process mining, as nodes only store their own as well as predecessor and successor events.
-To determine an event’s predecessor event, the node that records this particular event potentially has to query all other nodes, which can lead to large communication overheads.
-In this paper, we show that activities of real-world processes statistically only have a few predecessors activities, which further only rarely change throughout the lifetime of a process.
-Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [13] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors.
-As a result, for each event, a node can very efficiently determine the predecessor reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability.
-Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events.
-Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
-In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor and a baseline that immediately queries the correct node.
-Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner and show that in four out of five datasets an intermediate result FM already complies with the FM of the Alpha Miner by more than 90% after less than 5% of the events of the datasets.
-EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [14].
+Thus, over time, each node constructs a partial FM that represents the view of this node on the direct successions of this process.
+To derive the complete FM of the process, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix.
+As a result, EdgeMiner enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with the requesting entity and (b) privacy-preserving process mining, as nodes locally merely store their own as well as predecessor and successor events and a requesting entity only receives aggregate FMs instead of raw event streams.
+We prove, in this paper, that EdgeMiner guarantees an FM identical to the FM when using a centralized approach.
+Analyzing real-world datasets, our evaluation shows that activities statistically only have a few predecessor activities, which further only rarely change throughout the lifetime of a process.
+For example, even in the large Hospital Log (van Dongen, 2011), comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors.
+As a result, for each event, a node can very efficiently determine the predecessor, reducing the communication overhead by up to 96% compared to the naive baseline of querying all nodes in the network.
+Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeMiner can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
+As a final result, we show that for most datasets, an intermediate FM already complies with the final FM by more than 90% after less than 5% of the events of the dataset.
 In summary, the contributions of this paper are as follows.
-  • We introduce EdgeAlpha, a distributed Alpha Miner operating directly where the events originate.
-  • We analytically and experimentally show the correctness of the algorithm.
-  • We demonstrate that EdgeAlpha’s intermediate results quickly converge to the final result of the Alpha Miner.
-  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach and by 97.5% when batching queries with a batch size of 40.
-  • We ensure that the nodes only share aggregated event data with a central entity.
-  • We implement EdgeAlpha and publish its code open-source [^1] Available at: redacted..
+  • We introduce EdgeMiner, a distributed process mining algorithm enabling distributed process discovery and conformance checking by operating directly where the events originate.
+  • We ensure that the nodes only share aggregated event data with a requesting entity.
+  • We analytically and experimentally show the correctness of the merge of EdgeMiner.
+  • We demonstrate that EdgeMiner’s interim results quickly converge to the footprint matrix of the whole event log.
+  • We show that EdgeMiner reduces the message count for the predecessor determination by up to 96% compared to the naive approach and, for example, by 97.5% when batching queries with a batch size of 40.
+  • We implement EdgeMiner and publish its code open-source [^1] Available at: https://github.com/ds-kiel/EdgeMiner.
 The remainder of the paper is structured as follows.
-Section II introduces the necessary background.
-Next, Section III presents related work on distributed process discovery.
-Afterwards, we introduce the design of EdgeAlpha in Section IV.
-In Section V, we evaluate EdgeAlpha.
-We conclude the paper in Section VI.
-II Preliminaries and Background
-In this section, we provide the required preliminaries and background.
-We provide definitions of direct successions and footprint matrices and recapitulate the steps of the Alpha Miner Algorithm [12] that are relevant to our approach.
-We refer to Peterson [15] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [16, 17].
-In the following, let $L$ be an event log over the activity set $T$.
-We define direct successions and the resulting relationships between two activities following the definitions by Gatta et al.
-[9].
-Definition 1 (Direct Succession)
+Section 2 introduces the necessary background.
+Next, Section 3 presents related work on distributed process mining.
+Afterwards, we introduce the design of EdgeMiner in Section 4.
+In Section 5, we evaluate EdgeMiner.
+We conclude the paper in Section 6.
+2. Preliminaries and Background
+In this section, we give a brief introduction into process mining and provide definitions of events, traces, event logs, direct successions and footprint matrices.
+In process mining, we commonly discover a process model from an event log (process discovery) or check that a process follows a given model (conformance checking) (van der Aalst, 2011; van der Aalst, 2016; van der Aalst et al., 2012; van der Aalst, 2012).
+A common way to represent a process model is a Petri net (Peterson, 1977; Murata, 1989).
+Events and Event Logs: Process mining algorithms operate on event logs.
+Each event in the log typically contains at least three attributes: an activity, a timestamp, and a case ID.
+The case ID is used to distinguish different process instances.
+A trace is a sequence of events corresponding to a particular case ordered by timestamp.
+Direct Successions: In the following, let $L$ be an event log over the activity set $T$.
+We define direct successions and therewith definable relationships between two activities, following the definition by Gatta et al.
+(Gatta et al., 2019).
+Definition 2.1 (Direct Succession).
 Activity $b$ directly succeeds $a$ in event log $L$, i.e.
 $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{0},\dots,t_{n-1})\in L$ and $j\in\{0,\dots,n-2\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
-Definition 2 (Causality, No Direct Succession)
-With the help of Definition 1 we now define causality (1) and no direct succession (2) between two activities.
-$a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (1)
-$a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
-Next, we recapitulate the initial steps of the Alpha Miner before later in the paper introducing a distributed variant.
-First, the algorithm identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$.
-Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively.
-Subsequently, the algorithm determines all direct successions in $L$ which can be stored in an FM.
-This FM is then the basis for the further calculations in the algorithm and the resulting Petri net.
-A FM stores information regarding the direct succession relationships between activities within the event log.
+Definition 2.2 (Causality, No Direct Succession).
+With the help of Definition 2.1 we now define causality (1) and no direct succession (2) between two activities.
+(1) $a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$
+(2) $a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$
+Footprint Matrix: A footprint matrix (FM) stores information regarding the direct succession relationships between activities within the event log.
 For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$.
-This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
-$FM_{ij}=1\quad$ $\Leftrightarrow\quad i>_{L}j$
+An entry denotes how often the activity corresponding to the row was followed by the activity of the column.
+This way, $FM_{ij}\in\mathbb{N}_{0}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
+$FM_{ij}>0\quad$ $\Leftrightarrow\quad i>_{L}j$
 $FM_{ij}=0\quad$ $\Leftrightarrow\quad i\not>_{L}j$
 For our design, we define partial footprint matrices containing only one specific column: For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
-III Related Work
-Van der Aalst [18] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log.
-Here, models rather than FMs are created in a distributed manner.
-Evermann and Assadipour [19] leverage Map-Reduce [20] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes.
+Alpha Miner: To illustrate process mining and its use of FMs, we recapitulate the main steps of the seminal process discovery algorithm Alpha Miner (Algorithm 1).
+Later, we discuss how newer discovery algorithms extend the concept of the alpha miner and its FMs.
+First, the algorithm denotes all activities occurring in the event log $L$ as $T_{L}$.
+Each activity of this set corresponds to a transition in the final output model computed by the algorithm.
+Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively.
+Subsequently, the algorithm determines all direct successions in $L$ which it stores in an FM.
+This FM is the basis for the further calculations in the algorithm and the resulting Petri net.
+Next, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets.
+Each element of the first set is directly followed by every element of the second set, although not vice versa.
+Moreover, activities within each set do not directly succeed one another.
+The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step.
+Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
+Algorithm 1 Alpha Miner
+1:$T_{L}=\{t\in T|\exists\sigma\in L:t\in\sigma\}$
+2:$T_{I}=\{first(\sigma)\in T|\forall\sigma\in L\}$
+3:$T_{O}=\{last(\sigma)\in T|\forall\sigma\in L\}$
+[…]
+[…]
+6:$P_{L}=\{p_{(A,B)}|(A,B)\in Y_{L}\}\cup\{i_{L},o_{L}\}$
+[…]
+8:$\alpha(L)=(P_{L},T_{L},F_{L})$
+To form the set $X_{L}$, we need to have knowledge of the $\rightarrow_{L}$ and #$\#_{L}$ relationships between activities.
+These relations solely depend on the direct successions between the activities.
+An FM stores exactly those direct successions.
+For the Alpha Miner, it is not necessary to know the number of direct successions, just whether they exist or not.
+Heuristics Miner and Inductive Miner: Newer process discovery algorithms, extent on the concept of the Alpha Miner while keeping the FM as a key building block.
+For example, the Heuristics Miner starts with the construction of a dependency graph (Weijters et al., 2006).
+To build the dependency graph, it calculates a so-called dependency measure between each two activities, depending on the number of direct successions between the activities.
+As an FM stores all direct successions, the miner use it for this calculation.
+The Inductive Miner also relies on direct successions stored in an FM.
+The core concept is to construct a directly-follows graph, which visually represents the FM along with start and end activity sets.
+From this graph, a process tree is formed by repeatedly partitioning the nodes of the directly-follows graph, capturing the structure of the process step by step.
+3. Related Work
+Today’s approaches to distributed process discovery focus primarily on multiple compute nodes to speed up the processing of ever-growing, centralized event logs: Van der Aalst (van der Aalst, 2013) presents one of the first methods for distributed process discovery.
+He introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log.
+In contrast to EdgeMiner, he creates models rather than FMs in a distributed manner.
+Others (Evermann and Assadipour, 2014) leverage Map-Reduce (Dean and Ghemawat, 2008) to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes to handle the ever-growing size of event logs.
 However, their approach still relies on an event log that is initially stored at a central location.
-Subsequent work [21] extends their approach to the Flexible Heuristics Miner [22] and evaluates the performance of the Map-Reduce-based Alpha Miner and the Flexible Heuristics Miner.
-While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only storing data aggregates at a central location.
+Later works  (Evermann, 2016) bring Map-Reduce to the Flexible Heuristics Miner (Weijters and Ribeiro, 2011).
 Gatta et al.
-[9] address privacy concerns in the healthcare domain: In their work, each hospital center aggregates its own event data into an FM, such that the others only see the aggregated versions of the data.
-Like other works [18, 19, 21], they subsequently merge the FMs.
-In contrast, EdgeAlpha distributes the FM computations to the sensor nodes that record the events, enabling data aggregation at an earlier stage.
-This allows organizations to collect and utilize data in aggregated form that would otherwise be restricted in its raw form due to data protection regulations.
-As a result of processing events directly at their sources, EdgeAlpha increases scalability and reduces privacy risks.
-IV Design
-In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
-IV-A Assumptions on Events and Event Logs
+(Gatta et al., 2019) extends the primary motivation of the previous work (i.e., performance and scalability) to include privacy concerns in the healthcare domain, but still focuses on one event log per hospital.
+In contrast to all these, EdgeMiner emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only sending aggregate information, i.e., partial FMs, to the requesting node.
+In the context of conference checking, the Single-Entry Single-Exit (SESE) breaks down large process models and event logs into smaller sub-processes to analyze them independently (Munoz-Gama et al., 2014).
+Similarly, passage-based decomposition (van der Aalst and Verbeek, 2014) and Projected Conformance Checking (PCC) (Leemans et al., 2018) split logs into sets of activities that share direct successor-predecessor relationships: By solving each set independently and merging the results, it distributes, for example, conformance checking across multiple nodes.
+Although these approaches efficiently handle large and complex event logs, they remain centralized.
+One recent work presents conformance checking in a distributed manner also based on footprint-matrices, yet – in contrast to EdgeMiner– their work does not cover the underlying mechanism to distribute the calculation of the matrices (Andersen et al., 2024).
+Streaming process mining (Burattin, 2022) takes a different approach to distributed process mining.
+It focuses mainly on the real-time analysis of streams of events (i.e., event streams).
+Generally, those algorithms maintain computational state according to on previous events (Burattin et al., 2014; Hassani et al., 2015; van Zelst and et.
+al, 2019; Burattin et al., 2012) and employ approximations such as lossy counting (Burattin et al., 2014).
+While those methods are also motivated by scalability and performance, they are not designed to compute FMs at the sensor nodes.
+In contrast, EdgeMiner distributes the FM computation and each sensor node computes a partial FM, enabling data aggregation at an earlier stage.
+As a result of processing events directly at their sources, EdgeMiner increases scalability, reduces privacy risks and enables both process discovery and online conformance checking.
+4. Design
+In this section, we begin by discussing our assumptions and setting.
+Next, we give a design overview of EdgeMiner, detail on its design, introduce optimizations, prove correctness, and discuss how EdgeMiner enables distribution of well-known process mining algorithms.
+4.1. Assumptions and Setting
+4.1.1. Assumptions on Events and Event Logs
 In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case.
-Our evaluation shows these assumptions are both realistic and applicable in numerous real-world scenarios.
-For example, in a hospital each patient is commonly assigned a fixed ID that can serve as a case ID and the events such as blood tests or ultrasound are predefined activities.
-Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log.
-Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log.
-As we operate on distributed event logs, we assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
-IV-B EdgeAlpha Setting
-First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing the EdgeAlpha algorithm.
-We employ EdgeAlpha in typical IoT application settings: A network of sensor nodes that are, for example, integrated into machines in a smart factory or smart health setting.
+We assume timestamps within the same case to be distinct, ensuring unique FMs.
+In our evaluation in Section 5, we show that our assumptions are both realistic and applicable in numerous real-world scenarios.
+4.1.2. EdgeMiner Setting
+EdgeMiner targets typical IoT application settings: A network of IoT nodes that are, for example, integrated into machines in a smart factory or smart health setting.
 These sensors can communicate with each other, either directly or via the Internet.
 Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications.
-Together, these allow a node to find the processor and successor of an event.
-Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP [23].
+Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP (Mills, 1985).
 This is essential, as accurate timestamps ensure proper event ordering.
@@
 Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
-IV-C EdgeAlpha Algorithm: Overview
-EdgeAlpha consists of two phases: Phase 1 is initiated upon event triggering, and Phase 2 is activated upon a process model request.
+4.2. EdgeMiner Algorithm
+EdgeMiner consists of two phases: Phase 1 operates continuously upon each event detected, and Phase 2 is executed upon a footprint matrix request.
 In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs.
 As a result, every node continuously updates its local, partial FM.
-Other than Phase 1, Phase 2 is activated on demand whenever an entity requests a process model.
-Then, all nodes send their partial FMs to a central location, which constructs an FM.
-We define both phases, starting with the basic version of Phase 1, which sends a message to all nodes in the system for each event, followed by a more efficient variant and event batching.
-IV-D Phase 1 – Event Ordering & Partial Footprint Matrices
-Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing.
+Other than Phase 1, Phase 2 runs on demand, whenever a (central) entity requests the FM.
+Then, all nodes send their partial FMs to the requesting entity, which constructs the FM.
+4.2.1. Phase 1 – Event Ordering & Partial Footprint Matrices
+Figure 2.
+Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner: Without a central entity, nodes determine the order of events collaboratively using message passing.
+In this example, event 1 is a start event.
+Therefore, after the node detects the event, it queries the other nodes for the predecessor event.
+In our example it does not get any positive responses, and, thus, denotes it detected a start event.
+Upon sensing event 2, the detecting node queries for the predecessor event, receives a response, listing event 1 as predecessor.
+It stores this information in its local FM.
+[…]
+[Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner ]Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeMiner: Without a central entity, nodes determine the order of events collaboratively using message passing.
 In this specific setting, event 1 is a start event.
@@
 It increments the value in the corresponding cell of its local FM.
+For simplicity, we begin by introducing a naive version of EdgeMiner, which is not communication efficient.
+Later, we introduce optimizations that ensure efficiency.
 When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig.
@@
 Subsequently, node $j$ saves $e$ as the successor event of its own event.
-If a node has to choose its predecessor from several events, it indicates communication delays which may cause incorrectly set predecessors.
-However, this is automatically rectified during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor.
-The node sets the correct successor and informs the node corresponding to the incorrect successor.
-This one starts the predecessor search again.
-The node rectifies the stored events once the correct event is identified and updates the counts in its partial FM accordingly.
-For this reason nodes store a partial FM that not only encodes whether a direct succession has occurred but also counts the occurrences.
-The rectification process works due to a single start event and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests.
-As soon as a node has more than one successor in a single trace the incorrectly set one is informed and an total order of events is established.
-Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM.
+Events which do not have successors for a long duration are flagged as end activities.
+If a node has to choose its predecessor from several potential predecessor events, this indicates that events are happening in quick succession and (a) communication is delayed or (b) multiple nodes are identifying predecessors in parallel.
+This may lead to temporarily incorrectly set predecessors.
+EdgeMiner automatically rectifies such inconsistencies during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor: The node sets the correct successor and informs the incorrect successor.
+Then, the incorrect one starts the predecessor search again.
+A node then corrects the counters in its local FM.
+The rectification process converges since each trace has a single start event, and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests.
+Formally, each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM.
 Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities.
@@
 If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event.
-A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
-EdgeAlpha’s design further incorporates a sliding-window-principle for discarding old data in order to respond to the requirements of the resource-restricted sensor nodes.
-With this principle, as soon as a certain memory size has been reached or a certain time has been exceeded, the node discards the oldest data when adding new data.
-However, the sliding window approach only gets beneficial with a vast amount of data.
-The event logs we consider in the evaluation do not fill enough memory to utilize the sliding window.
-IV-E Phase 2 – Requesting Process Data
-This phase begins when an outside entity, for example, a central node, requests a process model.
-During this phase, we request the local FMs of all nodes and start or end events within any case, see Fig.
+In EdgeMiner, a node searches for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
+4.2.2. Phase 2 – Requesting a Footprint Matrix
+This phase begins when an outside entity requests the FM.
+During this phase, the outside entity requests the local FMs of all nodes and start and end events within any case, see Fig.
 3.
 Subsequently, each node responds with its local FM.
-Before sending, the node changes all entries greater than 0 to 1, aggregating the data even more.
-This converts the local partial FM back into the conventional FM structure.
-Each node checks if events without a successor were recorded to determine whether its respective activity is an end activity.
-The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID.
-Further, we merge the start and end activities into two sets $T_{I}$ and $T_{O}$, following the conventions of the Alpha Miner.
-Next, the requesting unit proceeds with the remainder of the Alpha Miner algorithm, starting after the calculation of the FM, here $FM^{all}$.
-Figure 3: Phase 2 – Requesting Process Data: We request partial FMs and start/end activity flags from all nodes.
-Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the process model.
-IV-F Most-Frequent-Predecessors (MFPs)
-In Phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log.
-To reduce communication overhead, we search for the predecessor by querying nodes in order of their likelihood of being the predecessor.
-Thus, when searching for an event’s predecessor, a node first contacts the most likely predecessor, then the second most likely one, and so on.
-For this, during runtime each node continuously records how often each predecessor occurs.
-As we show later in a dataset analysis, most activities in real-world scenarios are preceded by a typically stable and small subset of activities, this approach prioritizes querying the most frequently occurring predecessors.
-However, when detecting a start event, the corresponding node still has to query every node in the network to ensure it is the start event.
-In our evaluation in Section V we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures scalability.
-IV-G Batching
-To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events.
+The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\mathbb{N}_{0}^{n\times n}$ by concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID.
+Next, this FM serves as a process model for conformance checking, which compares the next events, traces, or sets of traces.
+Further, process discovery algorithms such as the Alpha Miner (van der Aalst et al., 2004), the Heuristics Miner (Weijters et al., 2006), or the Inductive Miner (Leemans et al., 2013) can use this assembled FM for process discovery.
+Please note that the outside entity only receives aggregate information, i.e., the partial FMs.
+Raw event data is only processed on the sensor nodes, ensuring both scalability and data privacy.
+Figure 3.
+Phase 2 – Requesting a Footprint Matrix: We request partial FMs and start/end activity flags from all nodes.
+Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the footprint matrix.
+[…]
+[Phase 2 – Requesting a Footprint Matrix]We request partial FMs and start/end activity flags from all nodes.
+Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the footprint matrix.
+4.3. Optimzaitons
+Next, we introduce three optimizations: Most-Frequent-Predecessors (MFPs), batching and a sliding window to enhance efficiency.
+4.3.1. Most-Frequent-Predecessors (MFPs)
+In Phase 1 of EdgeMiner, a node – upon detecting an event – sends out requests to all other nodes to determine the event’s predecessor.
+This results in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log.
+To reduce communication overhead, EdgeMiner contacts nodes in order of their likelihood of being the predecessor, i.e., it begins with the node that in the past was the most frequent predecessor according to its local FM.
+We show in a dataset analysis in Section 5 that activities of real-world processes statistically only have a few predecessor activities, which further only rarely change throughout the lifetime of a process.
+For example, even for datasets with a substantial number of activities, such as the Hospital Log (van Dongen, 2011) comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors.
+As a result, for each event, a node can very efficiently determine the predecessor, reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability.
+Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events.
+4.3.2. Batching
+To further minimize communication overhead in EdgeMiner, we batch querying of predecessor events.
 Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message.
@@
 The node ceases to query additional nodes once the batch is empty.
-IV-H Correctness
-After introducing the algorithmic design of EdgeAlpha, we next prove its correctness.
-For this, we show in the following that EdgeAlpha outputs the same process models as the traditional centralized Alpha Miner algorithm.
+For example, by batching predecessor queries in batches of 40 events, we show that EdgeMiner can reduce the average number of queried nodes per event to less than 2.5% of all nodes.
+4.3.3. Sliding Window
+To reduce resource demands, EdgeMiner utilizes a sliding window: Once the local event log reaches a given memory size or information age, the node discards it.
+4.4. Correctness
+After introducing the algorithmic design of EdgeMiner, we next show its correctness.
+We claim that merging the partial FMs in EdgeMiner outputs the same FM as when calculating the FM based on a full event log.
+Here, we discuss the intuition of the proof, a formal proof is in Appendix A.
+By definition of the FM, there is one column and row per activity.
+An entry of the FM describes how often the activity of the row was followed by the activity of the column.
+A column of an FM therefore contains all predecessor relations of a specific activity.
+In EdgeMiner, the calculation of these columns is distributed to the nodes, and they are updated whenever the predecessor of an activity is found.
+Every predecessor relation of a node, and therefore activity, is found by the design of EdgeMiner and no direct successions are added to the FMs of multiple nodes but exactly one.
+Thus, locally created columns in concatenated form ordered by activity ID must correspond exactly to the FM that is obtained when the entire log is considered and all direct successions are entered into a large matrix.
+In our experimental evaluation in Section 5, we show (a) that both EdgeMiner and established central algorithms result in identical FMs and thereby process models and (b) that in four out of five datasets an intermediate FM already complies with the FM of the whole event log by more than 90% after less than 5% of the events of the datasets.
+4.5. Distributed Mining with EdgeMiner
+By distributing the computation of FMs, we can efficiently implement key process mining algorithms, such as, for example, the Alpha Miner, Heuristics Miner, and Inductive Miner, while also supporting online conformance checking in a scalable and privacy-preserving manner.
+Below, we outline how EdgeMiner distributes each of these algorithms.
+4.5.1. Distributed Alpha Miner
+The Alpha Miner relies on identifying direct successions in event logs (cf.
+Section 2).
+Hence, it builds on causal relations between activities.
+We leverage EdgeMiner to distribute the computation of the FM from which the Alpha Miner constructs the process model.
+The Alpha Miner expects an FM consisting only of ones and zeros, i.e., if there is directly-follows relation or not.
+As EdgeMiner counts the number of successions, we simply set FM entries greater than zero to one.
+4.5.2. Distributed Heuristics Miner
+In contrast to the Alpha Miner, the Heuristics Miner not only considers whether a direct succession is present, but also considers the frequency and dependency measure of activity relationships.
+With a request, EdgeMiner merges all partial FMs to a global FM at the requesting entity, which includes the frequencies of the direct successions.
+From these, it calculates the dependency measures of the activity relationships.
+Next, the Heuristics Miner constructs a process model based on these frequencies and dependency measures.
+4.5.3. Distributed Inductive Miner
+The Inductive Miner first constructs a directly-follows graph and then recursively splits it into subprocesses to derive a process tree as process model.
+To obtain the directly-follows graph the Inductive Miner needs the direct successions present in the event log as well as the set of start activities and the set of end activities.
+EdgeMiner offers the direct successions in the form of an FM along with the two sets.
+Just as for the Alpha Miner, the frequencies of the direct successions are not relevant, so that all entries greater than zero are set to one.
+4.5.4. Distributed Footprint-Based Online Conformance Checking
+Online conformance checking continuously monitors process executions and compares them against a reference model in real time.
+Using EdgeMiner, each node generates and maintains a local, partial FM that captures the current state of the process at that node.
+As new events are detected, the local FMs are updated, and each node can use its FM to perform conformance checking against a predefined process model.
+5. Evaluation
+In this section, we evaluate EdgeMiner in terms of correctness, accuracy, and performance.
+We begin by discussing implementation, baselines and datasets in Section 5.1.
+Next, Section 5.2 analyzes these datasets.
+In Section 5.3, we experimentally evaluate the correctness of EdgeMiner.
+We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section 5.4.
+Section 5.5 estimates FM sizes during an algorithm run, and Section 5.6 reflects on the evaluation results.
+5.1. Evaluation Setup
+5.1.1. Implementation
+We implement EdgeMiner in Python and use a network of docker containers to simulate the distributed nature of EdgeMiner.
+5.1.2. Baselines
+In our experiments, we compare EdgeMiner with and without MFP Requesting to two baselines: (1) ”Query All” represents the upper bound where all nodes are queried, while (2) a theoretical oracle-like approach serves as the lower bound.
+The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
+Further, we compare EdgeMiner to traditional process discovery algorithms (van der Aalst et al., 2004; Weijters et al., 2006; Leemans et al., 2013) to validate that the resulting FMs are identical.
+5.1.3. Datasets
+We conduct experiments on the five rel-world datasets (event logs): Hospital Log (van Dongen, 2011), Sepsis Cases (Mannhardt, 2016) and an IoT Event Log for Process Mining in Smart Factories [^2] We use the MainProcess.xes file of the non-cleaned version.
+(Malburg et al., 2022).
+These process-mining event logs naturally support geographical distribution over multiple nodes.
+To further substantiate our evaluation, we also include the BPI Challenge 2017 (van Dongen, 2017) and the Road Traffic Fine Management Process (de Leoni and Mannhardt, 2015) datasets.
+In our setting, each activity in a dataset represents a node in an IoT network.
+Table 1. Dataset properties and EdgeMiner performance (Standard deviation in parentheses if not stated otherwise).
+[…]
+Traffic Fine 561,470 11 (1) 150,370 3.73 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
+Sepsis 15,214 16 (6) 1,050 14.49 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
+Smart Factories 8,607 21 (1) 271 31.76 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
+BPIC 2017 1,202,267 26 (1) 31,509 38.16 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
+Hospital 150,291 624 (29) 1,143 131.49 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
+5.2. Dataset Analysis
+We begin our experimental evaluation with an analysis of our real-world event logs, see Table 1.
+Our analysis shows that even in large datasets like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event.
+Further, our analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events.
+This characteristic benefits EdgeMiner, as self-loops do not require communication.
+5.3. Experimental Correctness & Fitness
+In addition to the analytical correctness proof in Section 4.4, we experimentally validate that the resulting FMs of EdgeMiner are identical to the ones computed by today’s centralized algorithms333we compare implementations in pm4py (https://pm4py.fit.fraunhofer.de).
+We test both on randomly generated artificial datasets and common real-world event logs, see Section 5.1.3.
+For all datasets the resulting FMs are identical at the end of the log, see Fig.
+7.
+Next, we examine how the merged FMs converge over time: Fig.
+7 demonstrates that nearly all datasets reach 90% fitness within less than 5% of the trace Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, as it is quite small, see Table 1.
+The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events before requesting the first FM.
+Moreover, our analysis suggests that reducing the sliding window for discarding old data has minimal impact on the resulting FM.
+5.4. Number of Queried Nodes
+To assess EdgeMiner’s performance, we analyze its two phases separately. For the first experiments, we assume a batch size of one.
+Figure 4.
+Average number of nodes queried with and without MFP Requesting, including standard error.
+MFP Requesting and knowledge of the start events reduce communication demands by a factor of 7.5 to 30 depending on the dataset.
+[…]
+[Comparison of average queried nodes with MFP Requesting and querying all nodes]Average number of nodes queried with and without MFP Requesting, including standard error.
+MFP Requesting and knowledge of the start events reduce communication demands by a factor of 7.5 to 30 depending on the dataset.
+Figure 5. Fitness over time. Intermediate FMs quickly converge to the centralized computed FM. The BPIC 2017 dataset, for example, already has a fitness of over 90% after 200 events.
+[…]
+[Fitness over time.]Intermediate FMs quickly converge to the centralized computed FM. The BPIC 2017 data set, for example, already has a fitness of over 90% after 200 events.
+Figure 6. Cumulative Distribution Functions: A steeper curve indicates a higher proportion of events that queried few nodes to identify their predecessors, reflecting greater efficiency.
+[…]
+[Cumulative Distribution Functions]Cumulative Distribution Functions: A steeper curve indicates a higher proportion of events that queried few nodes to identify their predecessors, reflecting greater efficiency.
+Figure 7. Relationship between batch size and average number of queried nodes per event, normalized. Batching efficiently reduces the number of queries.
+[…]
+[Relationship between batch size and average number of queried nodes per event]Relationship between batch size and average number of queried nodes per event, normalized.
+Batching efficiently reduces the number of quries.
+Figure 8.
+Number of nodes queried over time in the Hospital Log (moving average; green line: window size 50, step 1; orange line: window size 200, step 100).
+The moving average stabilizes after around 3,000 events.
+A reference line (Query All) indicates the maximum number of requested nodes per event.
+Note that the plots are limited to the first 10,000 of over 150,000 events.
+[…]
+[Number of nodes queried over time in the Hospital Log]Number of nodes queried over time in the Hospital Log (moving average; green line: window size 50, step 1; orange line: window size 200, step 100).
+The moving average stabilizes after around 3,000 events.
+A reference line (Query All) indicates the maximum number of requested nodes per event.
+Note that the plots are limited to the first 10,000 of over 150,000 events.
+Figure 9.
+Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis).
+The x-axis displays the different activities of the event log and highlights the start activities (S).
+We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all events using MFP Requesting.
+The diagram shows that activities with the highest average queries occur the least, excluding start activities.
+[…]
+[Bar diagram based on Sepsis Cases]Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis).
+The x-axis displays the different activities of the event log and highlights the start activities (S).
+The diagram shows that activities with the highest average queries occur the least, excluding start activities.
+5.4.1. Comparison Across Datasets
+Fig.
+4 and Table 1 illustrate, for each dataset, the average number of nodes queried per event to find the predecessor.
+Query All serves as the naive baseline of messaging every node in the network.
+With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event.
+We achieve additional message reduction when the process has a designated start event, making predecessor requests for the respective node redundant.
+Overall, the reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
+Fig.
+7 shows Cumulative Distribution Functions (CDF) for all datasets.
+The Hospital Log has the steepest increase, with over 90% of events needing queries to no more than 5% of the nodes.
+Conversely, the Traffic Fine dataset has very few instances where the predecessor can be identified after querying just 5% of the nodes.
+This deviation stems from the limited number of 11 activities in the Traffic Fine dataset.
+Hence, within the 5% threshold, only self-loops are considered.
+5.4.2. Stabilization over Time
+We now delve deeper into the specifics of the Hospital Log, as it is by far the largest one in terms of activities.
+Fig.
+8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events.
+There are still fluctuations, for example, at around 5,000 events.
+This event log is a good example of stabilization due to its large activity set.
+Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
+5.4.3. Activity-wise Analysis
+Next, we evaluate the number of queries per activity in the Sepsis Cases log, see Fig.
+9.
+In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones.
+However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others.
+Although ”Release E” shows an increased query count, it is not a start activity.
+The increased query count results from the limited occurrence (only six times), indicating non converged predecessor stabilization.
+The situation is similar for the four activities to its left.
+They occur infrequently, which also explains the higher standard error.
+5.4.4. Batching
+Fig.
+7 illustrates the reduction in the average number of queried nodes per event as the batch size increases.
+For example, at a batch size of 40, all datasets achieve an average number of nodes queried per event of less than 2.5% of the nodes in the network.
+For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
+5.4.5. Communication Cost and Privacy
+Table 2 compares the centralized approach with the naive EdgeMiner algorithm and EdgeMiner using MFP Requesting and batching in terms of communication overhead and privacy.
+The centralized approach requires fewer messages per event, but at the cost of reduced privacy.
+For the naive EdgeMiner, the number of messages per event is equal to $12\cdot(n-1)+$ 1 with $n$ the number of activities/nodes.
+All nodes except the node itself are contacted, all other nodes reply, and the chosen node gets a confirmation message.
+For EdgeMiner with MFP Requesting and batching we compute $12\cdot k+$ 1 with $k$ the average number of queried nodes per event (see Fig.
+4).
+In the centralized approach, event data is sent to a central entity right after the event, which requires only one message per event.
+However, this approach inherently relies on a central storage system where all events are collected, presenting privacy and scalability risks.
+Although the number of messages per event for EdgeMiner with MFP Requesting and batching is higher than in the centralized approach, it is a considerable improvement compared to the naive EdgeMiner.
+We use a batch size of 10 events.
+As we show in Section 5.4.4, batching with larger batch sizes reduces the number of queried nodes and therefore the number of messages per event even further.
+The trade-off presented here is clear: while the EdgeMiner variants, particularly EdgeMiner with MFP Requesting and batching, cause higher communication overhead, they offer substantial scalability and privacy benefits by eliminating central storage and limiting non-aggregated communication to within the network.
+Table 2. EdgeMiner’s efficiency and privacy across datasets against Centralized PM (CS/CE = Central Storage/Entity).
+Centr. Naive EdgeMiner EdgeMiner with MFP & Batching (10)
+Traffic Fine 1 21 1.76
+Sepsis 1 31 2.32
+Smart Factories 1 41 2.30
+BPIC 2017 1 51 1.76
+# Msg. per Event Hospital 1 1247 26.46
+Storage None No CS No CS
+Privacy Enablers Msg. Scope To CE In-Network In-Network
+5.5. Footprint Matrix Sizes
+In the second phase of EdgeMiner, nodes transmit their data to a (central) entity, generating a fixed number of messages.
+These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace.
+In EdgeMiner, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities.
+Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements.
+Compared to existing distributed process discovery algorithms (Evermann and Assadipour, 2014; Gatta et al., 2019), which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size.
+Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources.
+For long-running processes, EdgeMiner can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
+5.6. Evaluation Discussion
+EdgeMiner leverages data locality to enhance its efficiency.
+Unlike existing distributed process discovery algorithms (Evermann and Assadipour, 2014; Gatta et al., 2019), which typically collect the event data completely before processing it and creating FMs or process models, EdgeMiner adopts a proactive strategy by processing data at runtime and at its source.
+When the FM is requested, EdgeMiner has already started the process discovery algorithm, by collecting partial FMs at each node, and thereby speeding up mining.
+Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events.
+When looking at conformance checking, EdgeMiner enables online conformance checking by distributing the computation of FMs and making the partial FMs available at the data sources.
+6. Conclusion
+In this paper, we introduce EdgeMiner, a novel approach tailored to enable distributed process discovery and conformance checking in IoT networks.
+By distributing computations to IoT and edge devices, EdgeMiner addresses scalability and privacy concerns inherent in centralized approaches.
+We discuss how EdgeMiner’s design principles support common process mining algorithms that construct footprint matrices or record direct successions, such as the Alpha, Heuristics and Inductive Miner (van der Aalst et al., 2004; Weijters et al., 2006; Leemans et al., 2013).
+Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network.
+The batching of predecessor queries further minimizes communication overhead.
+Our empirical evaluations showcase EdgeMiner’s stark reduction of messages in real-world scenarios compared to a naive predecessor determination approach.
+EdgeMiner achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event.
+By batching predecessor queries in groups of, for example, 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes.
+We analytically and experimentally demonstrate the correctness of the algorithm and show how the intermediate results of EdgeMiner quickly converge to the FM of the full event log.
+We highlight how EdgeMiner provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same FMs as the centralized algorithms.
+By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeMiner enables data privacy through aggregation that eliminates personal identification.
+Acknowledgements.
+This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.
+Appendix A Appendix: Formal Correctness Proof
+We introduce a formal correctness proof. For this, we show that the merging of partial FMs in EdgeMiner outputs the same FM as calculating the FM based on a full event log.
 We begin by introducing notations before defining our claim.
-Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes all activities in $L$.
+Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes the set of all activities in $L$.
 Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of.
 Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column.
-In the following correctness claim, we omit start and end activities.
-The design of EdgeAlpha ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace.
-By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
-Claim 1
+In the following correctness claim, we only consider the footprint matrices and omit start and end activities.
+The design of EdgeMiner ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace.
+Further, we assume that the footprint matrices only consist of ones and zeroes.
+The correctness for footprint matrices with entries in $\mathbb{N}_{0}$ follows from the fact that EdgeMiner detects all direct successions and adds a direct succession to exactly one partial footprint matrix.
+Claim 1.
 Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity ID results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
-Proof 1
+Proof.
 To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa.
@@
 Thus proving the claim.
-$\Box$
-IV-I Design Discussion
-A major advantage of EdgeAlpha is that event logs are no longer stored in one central place.
-Instead each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity.
-In addition, we use batching, i.e.
-a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
-Further, the design of EdgeAlpha directly applies to process mining algorithms beyond the Alpha Miner.
-Algorithms that start by building an FM or keeping track of direct successions in the event log can be projected onto our distribution approach.
-A well-known example of another algorithm is the Heuristics Miner [14].
-The vast majority of results that we show in our evaluation in Section V, such as the efficiency of Phase 1 of EdgeAlpha, are not specific to the Alpha Miner and directly apply, for example, to the Heuristics Miner.
-V Evaluation
-In this section, we evaluate EdgeAlpha in terms of correctness, accuracy, and performance.
-We present our implementation in Section V-A and suitable baselines in Section V-B.
-Section V-C introduces on which well-known datasets we conduct the experiments.
-We analyze these datasets in Section V-D.
-In Section V-E we discuss correctness experiments and the compliance between the FMs of EdgeAlpha and the traditional Alpha Miner.
-We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section V-F.
-Section V-G estimates FM sizes during an algorithm run and Section V-H discusses the overall scalability of EdgeAlpha.
-V-A Implementation
-We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha.
-V-B Baselines
-In our experiments, we compare EdgeAlpha with and without MFP Requesting.
-”Query All” represents the upper bound where all nodes are queried, while a theoretical oracle-like approach serves as the lower bound.
-The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
-EdgeAlpha’s design assumes event detection as part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning.
-Thus, comparing EdgeAlpha to the traditional Alpha Miner [12] or other distributed Alpha Miners such as the Map-Reduce approach [19] or the Alpha Miner in the healthcare scenario [9] is not viable.
-The novelty of EdgeAlpha lies in utilizing the event detection phase to pre-process events, a feature absent in other approaches.
-V-C Datasets
-We conduct experiments on the datasets Hospital Log [13], Sepsis Cases [24] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version.
-[3].
-These datasets naturally support geographical distribution over multiple nodes.
-To further substantiate our evaluation, we also include the BPI Challenge 2017 [25] and the Road Traffic Fine Management Process [26].
-In our setting, each activity in a dataset represents a node in an IoT network.
-TABLE I: Dataset Characteristics (Standard deviation in parentheses if not stated otherwise).
-[…]
-Traffic Fine 561,470 11 (1) 150,370 3.73 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
-Sepsis 15,214 16 (6) 1,050 14.49 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
-Smart Factories 8,607 21 (1) 271 31.76 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
-BPIC 2017 1,202,267 26 (1) 31,509 38.16 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
-Hospital 150,291 624 (29) 1,143 131.49 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
-V-D Dataset Analysis
-We analyze the properties of real-world event logs.
-Next to basic information such as the numbers of events, activities and cases and the average case length, Table I includes data that directly influences the number of requests in Phase 1.
-These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity.
-Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
-Even in a dataset like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event.
-The dataset analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events.
-This characteristic benefits EdgeAlpha, as self-loops do not require communication.
-For self-loops, too, only in case of communication delays a self-loop may be set wrongfully for which, however, EdgeAlpha provides a rectification procedure.
-V-E Experimental Correctness and Fitness Analysis
-As a first step, we validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^3] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section V-C.
-Next, we examine how the merged FMs obtained by querying the nodes for their partial FMs after an increasing number of events differ from the intermediate FM produced by the traditional Alpha Miner.
-We compare them by using the footprint-based fitness value implemented in pm4py.
-Fig.
-7 confirms the full compliance of EdgeAlpha’s resulting FM with the Alpha Miner’s FM across all dataset.
-It also demonstrates that nearly all datasets reach 90% fitness with less than 5% of events.
-Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, although it is considerably small, as shown in Table I.
-The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events there are typically present in a common event log before requesting the first process model.
-Moreover, our analysis suggests that reducing the sliding window for discarding old data would have minimal impact on the resulting process model, but is dependent on the current dataset.
-V-F Number of Queried Nodes
-To assess EdgeAlpha’s performance, we analyze its two phases separately.
-The event ordering phase depends highly on the number of potential predecessors a node queries, leading to significant network traffic, particularly during concurrent events, in datasets like the Hospital Log, where up to 623 nodes are queried.
-MFP Requesting in EdgeAlpha builds on past event data to optimize requests in the future.
-For the first experiments, we assume a batch size of one to focus on other characteristics.
-As our batch size analysis shows, the different batch sizes strongly influence the number of queried nodes.
-Figure 4: Comparison of average queried nodes with MFP Requesting and querying all nodes.
-Diagram also shows standard error.
-EdgeAlpha can reduce communication by querying the most frequent predecessors.
-With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
-Figure 5: Fitness over time.
-Intermediate result FMs of the data sets quickly converge into the FM of the traditional Alpha Miner.
-The BPIC 2017 data set, for example, already has a fitness value of over 90% after 200 events compared with the Alpha Miner FM, which is approx.
-0.02% of the events in the event log.
-Figure 6: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence.
-A steeper curve indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
-Figure 7: Relationship between batch size and average number of queried nodes per event.
-Normalizing the average queried nodes for comparison.
-Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50.
-With increasing batch size, the reduction rate decreases.
-Figure 8: Stabilization of queried node count over time in Hospital Log.
-Since the data fluctuates, we use moving averages to visualize the stabilization.
-A reference line (Query All) indicates the maximum number of requested nodes per event.
-The moving average stabilizes after fewer than 3,000 events.
-We plot an additional line with a larger window size to show the trend of the stabilization.
-The plots are limited to the first 6,000 events.
-Figure 9: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis).
-The x-axis displays the different activities of the event log and highlights the start activities (S).
-We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all event.
-The diagram shows that activities with the highest average queries occur the least, excluding start activities.
-V-F1 Comparison Across Datasets
-Fig.
-4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor.
-Query All denotes the naive way to message every node in the network.
-This serves as a baseline to compare to.
-With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event.
-We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant.
-Table I shows the significant reduction in the number of queried nodes.
-This reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
-Fig.
-7 shows Cumulative Distribution Functions (CDF) for all datasets.
-The Hospital Log has the steepest increase with over 90% of events needing queries to no more than 5% of the nodes.
-Conversely, the Traffic Fine dataset has few instances where the predecessor can be identified after querying just 5% of the nodes.
-This deviation stems from the limited number of 11 activities in the Traffic Fine dataset.
-Hence, within the 5% threshold, only self-loops are considered.
-V-F2 Stabilization Analysis
-We now delve deeper into the specifics of the Hospital Log.
-Fig.
-8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events.
-There is still a higher amplitude after 5,000 events.
-However, there are no long-lasting deviations.
-This event log is a good example of stabilization due to its large activity set.
-Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
-V-F3 Activity-wise Analysis
-Fig.
-9 takes a closer look at the specifics of the Sepsis Cases log.
-Depending on the activity, the Sepsis Cases log shows varied querying patterns.
-In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones.
-However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others.
-Although ”Release E” shows an increased query count, it is not a start activity.
-The increased query count results from the limited occurrence (only six times), indicating ongoing stabilization.
-The situation is similar for the four activities to its left.
-They occur infrequently, which also explains the higher standard error.
-V-F4 Batch Size Analysis
-Fig.
-7 illustrates the reduction in the average number of queried nodes per event as the batch size increases.
-The rate in which the number of queries decreases varies between the datasets.
-However, for all datasets the rate decreases with a growing batch size.
-For all datasets, the average number of queried nodes per event already lies under 2.5% of nodes in the network with a batch size of 40.
-For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
-Overall, our results indicate that the reduction in the number of queries depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors per activity.
-The experiments show that it takes the nodes some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes.
-While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes.
-One such adaptation involves employing a smaller sliding window, where nodes track activity predecessors only within a fixed window.
-This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
-V-G Footprint Matrix Sizes
-In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages.
-These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace.
-Thus, the scalability of these messages depends highly on the size of each partial FM.
-In EdgeAlpha, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities.
-Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements.
-Compared to existing distributed Alpha Miners [19, 9], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size.
-Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources.
-For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
-V-H Evaluation Discussion
-EdgeAlpha leverages data locality to enhance its efficiency.
-Unlike existing distributed Alpha Miners [19, 9], which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source.
-As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm.
-Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events.
-Thus, building a process model becomes substantially faster.
-This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
-VI Conclusion
-In this paper, we introduce EdgeAlpha, a novel distributed approach tailored to process discovery in IoT networks.
-By distributing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches.
-We discuss how EdgeAlpha’s design principles can be applied to other process mining algorithms that initially construct footprint matrices or record direct successions, such as the Heuristics Miner [14].
-Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network.
-The batching of predecessor queries further minimizes communication overhead.
-Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach.
-EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event.
-By batching predecessor queries in groups of 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes.
-We analytically and experimentally demonstrate the correctness of the algorithm and additionally show how the intermediate results of EdgeAlpha quickly converge to the final result of the Alpha Miner.
-We highlight how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same results as the Alpha Miner.
-By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.
+∎